OpenStudy (anonymous):

What is 9th term of the arithmetic sequence whose 2nd and 6th terms are 107 and 87, respectively?

OpenStudy (anonymous):

@jakashaka123

OpenStudy (anonymous):

17?

OpenStudy (anonymous):

To find the common difference d, use the formula a1+4d=a5. For us, a1 is 10 and a5 is 38. 10+4d=38 Now we can solve for d. 4d=28 d=7 Add the common difference to the first term to get the second term. a2=a1+d=10+7=17

OpenStudy (anonymous):

yes its 17

OpenStudy (anonymous):

Otay :)

OpenStudy (anonymous):

np

OpenStudy (anonymous):

No its not 17 :P

OpenStudy (anonymous):

Mike .-. Stop bursting my bubble D':<

OpenStudy (anonymous):

I know trust me, ive literally been up 12 hours trying to do this lol. ive been wrong almost all night xD

OpenStudy (anonymous):

^^^ are the possible answers.

OpenStudy (anonymous):

Go slap that h o e they be wrong den!

OpenStudy (anonymous):

Lmfao

OpenStudy (anonymous):

▄︻̷̿┻̿═━一 The H O E ===> ʕ•ᴥ•ʔ

OpenStudy (anonymous):

(ง'̀-'́)ง

OpenStudy (anonymous):

Yeah I wish... I dont even know who the hoe is xD

OpenStudy (anonymous):

His name is sam! ┬┴┬┴┤(・_├┬┴┬┴ S H I T he's coming hide! I gtg ttyl

OpenStudy (anonymous):

Lol xD

OpenStudy (anonymous):

@jakashaka123 Can you help?

OpenStudy (anonymous):

@iGreen @ZeHanz

OpenStudy (anonymous):

@Vianne can you help?

OpenStudy (anonymous):

Any help would be great!

OpenStudy (zehanz):

There is no need for formulas in this case, because the numbers are small as are the differences between the terms involved. There are four steps to go from the second to the 6th term. each step is the same size (arithmetic seq). In those four steps you go from 107 to 87, so 20 less. That is 20/4=5 less each step. I think you know now how to find the ninth number :)

OpenStudy (anonymous):

so 122 then?

OpenStudy (zehanz):

No, because the sequence decreases: 2th term is 107, 6th is 87. Each next term is 5 less than the current one.

OpenStudy (anonymous):

ops sorry, wrong way, so 72 lmao

OpenStudy (zehanz):

Done!

OpenStudy (anonymous):

Thank you sooo much :D